# A Review of Bayesian Uncertainty Quantification in Deep Probabilistic Image Segmentation

### v1

@@ IV-A2 Spatial correlation
. To explicitly model spatial correlation within pixel space of the likelihood distribution, Monteiro et al. [50] propose the Stochastic Segmentation Network (SSN), which models the output logits as a multivariate normal distribution parameterized by the neural networks $f_{\bm{\theta}}^{\mu}$ and $f_{\bm{\theta}}^{\Sigma}$ as
@@
where the covariance matrix has low-rank structure
@@
Here, $\mathbf{P}$ has dimensionality $((H\times W\times K)\times R)$, with $R$ being a hyperparameter that controls the parameterization rank and $\bm{\Lambda}$ a diagonal matrix. This results in a more structured and expressive distribution, while retaining reasonable efficiency. As the SoftMax transform on this low-rank multivariate normal distribution pertains an intractable integral, Monte Carlo sampling is employed. SSNs can be augmented as an additional layer to any existing CNN (see Figure 5).
@@
The PixelCNN remains a popular method to model such relationship due to its substantial receptive field [51, 52]. Zhang et al. [53] suggest to use this to predict a downsampled segmentation mask $\mathbf{\tilde{y}}$ with $p_{\bm{\phi}}(\mathbf{\tilde{y}}|\mathbf{x}$), and fuse this with a conventional CNN to predict the full resolution mask with $p_{\bm{\theta}}(\mathbf{y}|\mathbf{\tilde{y}},\mathbf{x}$). Fusing the two masks is done through a resampling module, containing a series of specific transformations to improve quality and diversity of the samples. See Figure 6 for an illustration. Notably, PixelCNNs employ a recursive sampling process, which also enables completion/inpainting of user-given inputs.
Figure 6: An illustration of the PixelCNN-based PixelSeg [53]. ‘C’ indicat es the concatenation module, ’R’ the resampling module and $\sigma$ the softmax activation. Partially transparent elements appear during test-time sampling.
@@ IV-B Latent-level sampling
Directly learning the conditional data distribution is a challenging task. Hence, generative models often employ a simpler latent (unobserved) variable $\mathbf{Z}\sim p_{\mathbf{Z}}$ with $\mathcal{Z}\in\mathbb{R}^{d}$, to then learn the approximate joint density $p(\mathbf{Y}|\mathbf{Z},\mathbf{X})$. The marginalized distribution is obtained through decomposition
@@
with parameters $\bm{\theta},\bm{\psi}$. Conditioning the latent density on the input images is not a necessity but usually preferred for smooth optimization trajectories [54]. As such, the spatial correlation is not explicitly modeled but rather induced through mapping the latent variables to segmentation masks. Notable architectures relevant to the context of this paper are Generative Adversarial Networks (Section IV-B1), Variational Autoencoders (Section IV-B2) and Denoising Diffusion Probabilistic Models (Section IV-B3).
IV-B1 Generative Adversarial Networks
A straightforward approach is to simply learn the decomposition in Equation (14) through sampling from an unconditional prior density
$p_{\mathbf{Z}}=\mathcal{N}(\bm{\mu}=\mathbf{0},\bm{\Sigma}=\bm{\sigma}\cdot\mathbf{I}),$ (15)
and mapping this to segmentation $\mathbf{Y}$ through a generator $G_{\bm{\phi}}:\mathcal{X}\times\mathcal{Z}\rightarrow\mathcal{Y}$. Goodfellow et al.  [55] show that this approach can notably enhanced through the incorporation of a discriminative function (the discriminator), denoted as $D_{\bm{\psi}}:\mathbb{R}^{C\times D}\rightarrow[0,1]$. In this way, $G_{\bm{\phi}}$ learns reconstruct realistic looking images, guided by the discriminative capabilities of $D_{\bm{\psi}}$, making sufficient resistance from $D_{\bm{\psi}}$ to $G_{\bm{\phi}}$ imperative. We can denote the cost of $G_{\bm{\phi}}$ in the GAN as the negative cost of $D_{\bm{\psi}}$ as
$J_{G}=-J_{D}=$ $\,\mathbb{E}_{p_{\mathcal{D}}}[\,\log D_{\bm{\psi}}(\mathbf{y})\,]$ (16)
@@ IV-B1 Generative Adversarial Networks
Figure 7: The Calibrated Adversarial Refinement network [56] based on the Generative Adversarial Network with additional loss terms.
While conditional GANs had been used for semantic segmentation before [57], Kassapis et al. [56] explicitly contextualized the architecture within aleatoric uncertainty quantification using their proposed Calibrated Adversarial Refinement (CAR) network (see Figure 7). The calibration network, $F_{\bm{\theta}}:\mathbb{R}^{C\times D}\rightarrow\mathbb{R}^{K\times D}$, initially provides a SoftMax activated prediction as $F_{\bm{\theta}}(\mathbf{x})=\mathbf{c}$, with (cross entropy) reconstruction loss
$\mathcal{L}_{rec}=-\mathbb{E}_{p_{\mathcal{D}}}[\log p_{\bm{\theta}}(\mathbf{c}|\mathbf{x})].$ (17)
Then, the conditional refinement network $G_{\bm{\theta}}$ uses $\mathbf{c}$ together with $\mathbf{X}$ and latent samples $\mathbf{Z}_{i}\sim p_{\mathbf{Z}}$ injected at multiple decomposition scales $i$, to predict various segmentation maps. Furthermore, the refinement network is subject to adversarial objective
$\mathcal{L}_{adv}=-\mathbb{E}_{p_{\mathcal{D}}}\mathbb{E}_{p_{\mathbf{Z}}}[\log D_{\bm{\psi}}(G_{\bm{\phi}}(F_{\bm{\theta}}(\mathbf{x}),\mathbf{z}),\mathbf{x}),\mathbf{x})],$ (18)
which is argued to elicit superior structural qualities compared to relying on cross-entropy alone. At the same time, the discriminator opposes the optimization with
@@
Finally, the average of the $N$ segmentation maps generated from $G_{\phi}$ are then compared against the initial prediction of $F_{\theta}$ through the calibration loss, which is the analytical KL-divergence between the two categorical densities denoted as
$\mathcal{L}_{cal}=\mathbb{E}_{p_{\mathcal{D}}}\operatorname{KL}[\,p_{\bm{\phi}}(\mathbf{y}|\mathbf{c},\mathbf{x})||p_{\bm{\theta}}(\mathbf{c}|\mathbf{x})\,].$ (20)
@@
$\mathcal{L}_{G}=\mathcal{L}_{adv}+\lambda\cdot\mathcal{L}_{cal},$ (21)
with hyperparameter $\lambda\geq 0$. The purpose of the calibration network is argued to be three-fold. Namely, it (1) sets a calibration target for $\mathcal{L}_{cal}$, (2) provides an alternate representation of $\mathbf{X}$ to $G_{\bm{\phi}}$, and (3) allows for sample-free aleatoric uncertainty quantification. The refinement network can be seen as modeling the spatial dependence across the pixels, which enables sampling coherent segmentation maps through latent variable $\mathbf{Z}$.
IV-B2 Variational Autoencoders
Techniques such as GANs rely on implicit distributions and are void of any notion of data likelihoods. An alternative approach estimates the Bayesian posterior w.r.t. the latent variables, $p(\mathbf{Z}|\mathbf{Y},\mathbf{X})$, with an approximation $q_{\theta}(\mathbf{Z}|\mathbf{Y},\mathbf{X})$ obtained my maximizing conditional Evidence Lower Bound (ELBO)
This is also known as Variational Inference (VI). The first term in Equation (22) represents the reconstruction cost of the decoder subject to the latent code $\mathbf{Z}$ and input image $\mathbf{X}$. The second term is the Kullback-Leibler (KL) divergence between the approximate posterior and prior density. As a consequence of the mean-field approximation, all involved densities are modeled by axis-aligned Normal densities and amortized through neural networks parameterized by $\phi$, $\theta$ and $\psi$. The predictive distribution after observing dataset $\mathcal{D}$ is then obtained as
[…]
Implementing the conditional ELBO in Equation (22) can be achieved through a VAE-like architecture [10]. A few additional design choices specific for uncertainty quantification for segmentation result in the The Probabilistic U-Net (PU-Net) [58]. Firstly, the latent variable $\mathbf{z}$ is only introduced at the final layers of a U-Net conditioned on $\mathbf{X}$, where the vector is up-scaled through tiling and then concatenated with the feature maps of the penultimate layer, which is followed by a sequence of 1$\times$1 convolutions. When involving conditional latent variables in this manner, it is expected that most of the semantic feature extraction and delineation is performed in the U-Net, while the information $\mathbf{Z}$ provides is almost exclusively regarding the segmentation variability. Therefore, relatively smaller values of $d$ are feasible than what is commonly used in image generation tasks.
Figure 8: The Probabilistic U-Net [58] based on a conditional Variational Autoencoder [10]. Latent samples are inserted at the final stages of a U-Net through a tiling operation.
Similar to related research on the VAE [59, 60, 61, 62, 63], much work has been dedicated to improving the PU-Net. For instance, investigation on improving VI with novel parameterization of the amortized densities also provide interesting insights into model behaviour [64, 65, 66, 67, 68, 69, 70]. Furthermore, extending the architecture to multiple decomposition hierarchies [71, 72], three-dimensional modalities [73, 74, 75, 76, 77] and conditioning on the annotator [78] also resulted in substantial performance gains.
Density parameterization
Augmenting a Normalizing Flow (NF) to the posterior density of a VAE is a commonly used tactic to improve its expressiveness [63]. This phenomena has also been successful for cVAE-like models such as the PU-Net [64, 66]. NFs are a class of generative models that utilize $k$ consecutive bijective transformations $f_{k}:\mathbb{R}^{D}\rightarrow\mathbb{R}^{D}$ as $\mathbf{f}=f_{K}\circ\ldots\circ f_{k}\circ\ldots\circ f_{1}$, to express exact log-likelihoods of arbitrarily complex distributions $\log p(\mathbf{x}|\mathbf{z})$. These are often denoted as $\log p(\mathbf{x})$ for simplicity and can be determined through
@@ IV-B2 Variational Autoencoders
Bhat et al. [67, 68] compare this approach with other parameterizations of the latent space including a mixture of Gaussians and low-rank approximation of the full covariance matrix. Valiuddin et al. [79] show that the latent space can converge to contain non-informative latent dimensions, undermining the capabilities of the latent-variable approach, generally referred to as mode or posterior collapse [80, 59]. Their proposition considers the alternative formulation of the ELBO
@@
$\hskip 70.0pt-I(\mathbf{Y},\mathbf{Z}|\mathbf{X}),$
which the novel objective maximizes the contribution of the (expected) mutual information between latent and output variables. This can be rewritten to the objective
@@
with $\operatorname{S}_{\epsilon}$ being the Sinkhorn divergence [81] and $\alpha,\beta$ being hyperparameters, which results in a more uniform latent space leading to increased model performance. Also, modeling the ELBO of the joint density has also been explored [69]. This formulation results in an additional reconstruction term and forces the latent variables to be more congruent with the data. Furthermore, constraining the latent space to be discrete has resulted in some improvements, where it is hypothesized that this counters the model collapse phenomena in latent space [70].
@@
Learning latent features over several hierarchical scales can provide expressive densities and interpretable features across various abstraction levels [82, 83, 84, 85, 86]. Such models commonly fall under hierarchical VAE umbrella term. Often, an additional Markov assumption of length $T$ is placed on the posterior as
@@
This objective is implemented in the Hierarchical PU-Net [71] (HPU-Net, depicted in Figure 9). Simply stated, the architecture learns a latent representation at multiple decomposition levels of the U-Net. Furthermore, residual connections in the convolutional layers are necessary to prevent degeneracy of uninformative latent variables with the KL-divergence rapidly approaching zero. For similar reasons, the Generalized ELBO with Constrained Optimization (GECO) objective was employed, which extends on Equation (29) as
$\mathcal{L}_{GECO}=\lambda$ $\cdot(\mathbb{E}_{q_{\theta}(\mathbf{Z}|\mathbf{Y},\mathbf{X})}[\,\log p_{\phi}(\mathbf{Y}|\mathbf{z},\mathbf{X})\,]-\kappa)$ (30)
@@
Hyperparameter $\lambda$ is the Lagrange multiplier update through the Exponential Moving Average of the reconstruction, which is constrained to reach target value $\kappa$ set beforehand to an appropriate value. Finally, online negative hard mining is used to only backpropagate 2% of the worst performing pixels, which are stochastically picked with the Gumbel-SoftMax trick [87, 88]. PHiSeg [72] takes a similar approach to the HPU-Net. However, instead of placing the residual connections in the convolutional layers, PHiSeg uses these between the latent vectors across decomposition. Furthermore, auxiliary outputs, or deep supervision, at each decomposition scale is used to enforce disentanglement across latent variables.
@@
Early methods for uncertainty quantification in medical imaging primarily utilized 2D slices from three-dimensional (3D) datasets, leading to potential loss of critical spatial information and subtle nuances often necessary for accurate delineation. This limitation has spurred extensive research into 3D probabilistic segmentation techniques with ELBO-based models, aiming to preserve the integrity of entire 3D structures. Initial works [76, 77] demonstrate that the PU-Net can be adapted by replacing all 2D operations with their 3D variants. Crucially, the fusion of the latent sample with 3D extracted features is achieved through a 3D tiling operation. Viviers et al. [75] additionally augment a Normalizing Flow to the posterior density, as described in Section IV-B2. Further enhancements to the architecture include the implementation of the 3D hierarchical PU-Net [74] or an updated feature network incorporating the attention mechanisms, a nested decoder, and different reconstruction loss components tailored to specific applications [73].
@@
It can be relevant to model the annotators independently in cases with consistent annotator-segmentation pairs in the dataset. This can be achieved by conditioning the learned densities on the annotator itself [78, 89]. For example, features of a U-Net can be combined with samples from annotator specific Gaussian distributed posterior distributions [89]. Considering the approach from Gao et al. [78], generating a segmentation mask is achieved by first sampling an annotator from a categorical prior distribution $\mathcal{C}(\pi_{k}(\mathbf{x}))$, governed by the image conditional parameters $\pi_{k}(\mathbf{x})$ with for the $k$-th annotator. Then, samples are taken from its corresponding prior density as $\mathbf{z}_{k}\sim p_{k}(\mathbf{z}_{k})$ to reconstruct a segmentation through image-conditional decoder as $\mathbf{y}=F(\mathbf{x},\mathbf{z}_{k})$. The parameters $\pi_{k}(\mathbf{z}_{k})$ can also be used to weigh the corresponding predictions to express the uncertainty in the prediction ensemble. Additionally, consistency between the model and ground-truth distribution is additionally enforced through an optimal transport loss between the set of predictions and labels.
IV-B3 Denoising Diffusion Probabilistic Models
Recent developments in generative modeling have resulted a family of models known as Denoising Diffusion Probabilistic Models (DDPMs) [90, 91, 92]. Such models are especially renowned for their expressive power by able to encapsulate large and diverse datasets. While several perspectives can be used to introduce the the DDPMs, we build upon the earlier discussed HVAE (Section IV-B2) to maintain cohesiveness with the overall manuscript. In specific, let us introduce three additional modifications to the HVAE. Firstly, the latent dimensionality is set equal to the data dimensions, i.e. $d\,$=$\,D$. As a consequence, redundant notation of $\mathbf{Z}$ is removed and $\mathbf{Y}$ is instead subscripted with $t\in\{1,...,T\}$, indicating the encoding depth, where $\mathbf{Y}_{0}$ is the initial segmentation mask. Secondly, the encoding process (or forward process) is predefined as a linear Gaussian model such that
$q(\mathbf{y}_{T}|\mathbf{y}_{0})=p(\mathbf{y}_{0})\prod^{T}_{t=1}q(\mathbf{y}_{t}|\mathbf{y}_{t-1})$ (31)
and

### v3

@@ IV-A2 Spatial correlation
To explicitly model spatial correlation within pixel space of the likelihood distribution, Monteiro et al. [50] propose the Stochastic Segmentation Network (SSN), which models the output logits as a multivariate normal distribution, parameterized by the neural networks $f_{\bm{\theta}}^{\mu}$ and $f_{\bm{\theta}}^{\Sigma}$ as
@@
where the covariance matrix has a low-rank structure
@@
Here, $\mathbf{P}$ has dimensionality $((K\times D)\times R)$, with $R$ being a hyperparameter that controls the parameterization rank and $\bm{\Lambda}$ a diagonal matrix. This results in a more structured and expressive distribution, while retaining reasonable efficiency. As the SoftMax transform on this low-rank multivariate normal distribution pertains an intractable integral, Monte Carlo sampling is employed. SSNs can be augmented as an additional layer to any existing CNN (see Figure 5).
@@
A popular method to model such relationship due to its substantial receptive field is known as the PixelCNN [51, 52]. Zhang et al. [53] suggest to use this network to predict a downsampled segmentation mask $\mathbf{\tilde{y}}$ with $p_{\bm{\phi}}(\mathbf{\tilde{y}}|\mathbf{x}$), and fuse this with a conventional CNN to predict the full-resolution mask with $p_{\bm{\theta}}(\mathbf{y}|\mathbf{\tilde{y}},\mathbf{x}$). The two masks are fused through a resampling module, containing a series of specific transformations to improve quality and diversity of the samples. Figure 6 illustrates this concept. Notably, PixelCNNs employ a recursive sampling process, which also enables completion/inpainting of user-given inputs.
Figure 6: Illustration of the PixelCNN-based PixelSeg [53]. Parameter ‘C’ indicates the concatenation module, ’R’ the resampling module and $\sigma$ the softmax activation. Dotted elements appear during test-time sampling.
@@ IV-B Latent-level sampling
Directly learning the conditional data distribution is a challenging task. Hence, generative models often employ a simpler latent (unobserved) variable $\mathbf{Z}\sim p_{\mathbf{Z}}$ with $\mathcal{z}\in\mathbb{R}^{d}$, to instead learn the approximate joint density $p(\mathbf{Y}|\mathbf{Z},\mathbf{X})$. The marginalized distribution is obtained through integral
@@
with parameters $\bm{\theta},\bm{\psi}$. Conditioning the latent density on the input images is not a necessity, but usually preferred for smooth optimization trajectories [54]. As such, the spatial correlation is not explicitly modeled, rather induced through mapping the latent variables to segmentation masks. Notable generative model architectures relevant to this section are Generative Adversarial Networks (Section IV-B1), Variational Autoencoders (Section IV-B2) and Denoising Diffusion Probabilistic Models (Section IV-B3).
IV-B1 Generative Adversarial Networks (GANs)
A straightforward approach is to simply learn the marginalization in Equation (13) through sampling from an unconditional prior density
$p_{\mathbf{Z}}=\mathcal{N}(\bm{\mu}=\mathbf{0},\bm{\Sigma}=\mathbf{I}),$ (14)
and mapping this to segmentation $\mathbf{Y}$ through a generator $G_{\bm{\phi}}:\mathcal{X}\times\mathcal{Z}\rightarrow\mathcal{Y}$. Goodfellow et al.  [55] show that this approach can be notably enhanced through the incorporation of a discriminative function (the discriminator), denoted as $D_{\bm{\psi}}:\mathbb{R}^{C\times D}\rightarrow[0,1]$. In this way, $G_{\bm{\phi}}$ learns to reconstruct realistically looking images through the discriminative capabilities of $D_{\bm{\psi}}$, making sufficient guidance from $D_{\bm{\psi}}$ to $G_{\bm{\phi}}$ imperative. We can denote the cost of $G_{\bm{\phi}}$ in the GAN as the negative cost of $D_{\bm{\psi}}$ as
$J_{G_{\bm{\phi}}}=-J_{D_{\bm{\psi}}}=$ $\,\mathbb{E}_{p_{\mathcal{D}}}[\,\log D_{\bm{\psi}}(\mathbf{y})\,]$ (15)
@@ IV-B1 Generative Adversarial Networks (GANs)
Figure 7: Diagram of the Calibrated Adversarial Refinement network [56], based on the Generative Adversarial Network with additional loss terms.
While conditional GANs have been used for semantic segmentation earlier [57], Kassapis et al. [56] explicitly contextualized the architecture within aleatoric uncertainty quantification, using their proposed Calibrated Adversarial Refinement (CAR) network (see Figure 7). The calibration network, $F_{\bm{\theta}}:\mathbb{R}^{C\times D}\rightarrow\mathbb{R}^{K\times D}$, initially provides a SoftMax activated prediction as $F_{\bm{\theta}}(\mathbf{x})=\mathbf{c}$, with (cross-entropy) reconstruction loss
$\mathcal{L}_{\text{rec}}=-\mathbb{E}_{p_{\mathcal{D}}}[\log p_{\bm{\theta}}(\mathbf{c}|\mathbf{x})].$ (16)
Then, the conditional refinement network $G_{\bm{\theta}}$ uses $\mathbf{c}$ together with $\mathbf{X}$ and latent samples $\mathbf{Z}_{i}\sim p_{\mathbf{Z}}$ injected at multiple decomposition scales $i$, to predict various segmentation maps. Furthermore, the refinement network is subject to the adversarial objective
$\mathcal{L}_{\text{adv}}=-\mathbb{E}_{p_{\mathcal{D}}}\mathbb{E}_{p_{\mathbf{Z}}}[\log D_{\bm{\psi}}(G_{\bm{\phi}}(F_{\bm{\theta}}(\mathbf{x}),\mathbf{z}),\mathbf{x}),\mathbf{x})],$ (17)
which is argued to elicit superior structural qualities compared to relying solely on cross-entropy loss. At the same time, the discriminator opposes the optimization with
@@
Finally, the average of the $N$ segmentation maps generated from $G_{\phi}$ are compared against the initial prediction of $F_{\theta}$ through the calibration loss, which is the analytical KL-divergence between the two categorical densities, denoted by
$\mathcal{L}_{\text{cal}}=\mathbb{E}_{p_{\mathcal{D}}}\operatorname{KL}[\,p_{\bm{\phi}}(\mathbf{y}|\mathbf{c},\mathbf{x})\,||\,p_{\bm{\theta}}(\mathbf{c}|\mathbf{x})\,].$ (19)
@@
$\mathcal{L}_{G}=\mathcal{L}_{\text{adv}}+\lambda\cdot\mathcal{L}_{\text{cal}},$ (20)
with hyperparameter $\lambda\geq 0$. The purpose of the calibration network is argued to be threefold. Namely, it (1) sets a calibration target for $\mathcal{L}_{\text{cal}}$, (2) provides an alternate representation of $\mathbf{X}$ to $G_{\bm{\phi}}$, and (3) allows for sample-free aleatoric uncertainty quantification. The refinement network can be seen as modeling the spatial dependency across the pixels, which enables sampling coherent segmentation maps through latent variable $\mathbf{Z}$.
IV-B2 Variational Autoencoders (VAEs)
Augmenting a Normalizing Flow (NF) to the posterior density of a VAE is a commonly used tactic to improve its expressiveness [58]. This phenomenon has also been successfully applied to VAE-like models such as the PU-Net [59, 60]. NFs are a class of generative models that utilize $k$ consecutive bijective transformations $f_{k}:\mathbb{R}^{D}\rightarrow\mathbb{R}^{D}$ as $\mathbf{f}=f_{K}\circ\ldots\circ f_{k}\circ\ldots\circ f_{1}$, to express exact log-likelihoods of arbitrarily complex distributions $\log p(\mathbf{x}|\mathbf{z})$, where symbol $\circ$ indicates functional composition. These are often denoted as $\log p(\mathbf{x})$ for simplicity and can be determined through
@@ IV-B2 Variational Autoencoders (VAEs)
Bhat et al. [61, 62] compare this approach with other parameterizations of the latent space, including a mixture of Gaussians and low-rank approximation of the full covariance matrix. Valiuddin et al. [63] show that the latent space can converge to contain non-informative latent dimensions, undermining the capabilities of the latent-variable approach, generally referred to as mode or posterior collapse [64, 65]. Their proposition considers the alternative formulation of the ELBO, specified as
@@
$\hskip 70.0pt-I(\mathbf{Y},\mathbf{Z}|\mathbf{X}).$
The proposed novel objective maximizes the contribution of the (expected) mutual information between latent and output variables. This can be rewritten into the objective
@@
with $\operatorname{S}_{\epsilon}$ being the Sinkhorn divergence [66] and $\alpha,\beta$ denoting hyperparameters. This adaptation results in a more uniform latent space leading to increased model performance. Also, modeling the ELBO of the joint density has been explored [67]. This formulation results in an additional reconstruction term and forces the latent variables to be more congruent with the data. Furthermore, constraining the latent space to be discrete has resulted in some improvements, where it is hypothesized that this partially addresses the model collapse phenomenon [68].
Density parameterization
Techniques such as GANs rely on implicit distributions and are void of any notion of data likelihoods. An alternative approach estimates the Bayesian posterior w.r.t. the latent variables, $p(\mathbf{Z}|\mathbf{Y},\mathbf{X})$, with an approximation $q_{\theta}(\mathbf{Z}|\mathbf{Y},\mathbf{X})$, obtained by maximizing the conditional Evidence Lower Bound (ELBO)
This approach of estimating $p(\mathbf{Z}|\mathbf{Y},\mathbf{X})$ is also known as Variational Inference (VI). The first term in Equation (25) represents the reconstruction cost of the decoder subject to the latent code $\mathbf{Z}$ and input image $\mathbf{X}$. The second term is the Kullback-Leibler (KL) divergence between the approximate posterior and prior density. As a consequence of the mean-field approximation, all involved densities are modeled by axis-aligned Gaussian densities and amortized through neural networks, parameterized by $\phi$, $\theta$ and $\psi$. The predictive distribution after observing dataset $\mathcal{D}$ is then obtained as
$p(\mathbf{Y}|\mathbf{x}^{*})=\int p_{\phi}(\mathbf{Y}|\mathbf{z},\mathbf{x}^{*})q_{\theta}(\mathbf{z}|\mathcal{D})\mathrm{d}\mathbf{z}.$ (26)
The implementation of the conditional ELBO in Equation (25) can be achieved through a VAE-like architecture [10]. A few additional design choices specific for uncertainty quantification in segmentation lead to the Probabilistic U-Net (PU-Net) [69]. Firstly, the latent variable $\mathbf{Z}$ is only introduced at the final layers of a U-Net conditioned on $\mathbf{X}$, where the vector is up-scaled through tiling and then concatenated with the feature maps of the penultimate layer, which is followed by a sequence of 1$\times$1 convolution for classification. When involving conditional latent variables in this manner, it is expected that most of the semantic feature extraction and delineation are performed in the U-Net, while the provided information by $\mathbf{Z}$ is almost exclusively related to the segmentation variability. Therefore, relatively smaller values of $d$ are feasible than what is commonly used in conventional image generation tasks.
Figure 8: Depiction of the Probabilistic U-Net [69] based on a conditional Variational Autoencoder [10]. The latent samples are inserted at the final stages of a U-Net through a tiling operation.
Similar to related research on the VAE [65, 70, 71, 72, 58], much work has been dedicated to improving the PU-Net. For instance, investigation on improving VI with novel parameterization of the amortized densities also provides insights into model behaviour [59, 73, 60, 61, 62, 67, 68] (Section IV-B2). Furthermore, extending the architecture to multiple decomposition hierarchies [74, 75] (Section IV-B2), three-dimensional modalities [76, 77, 78, 79, 80] (Section IV-B2), and conditioning on the annotator [81] (Section IV-B2), also result in substantial performance gains.
@@
Learning latent features over several hierarchical scales can provide expressive densities and interpretable features across various abstraction levels [82, 83, 84, 85, 86]. Such models are commonly categorized under hierarchical VAE (HVAE) modeling. Often, an additional Markov assumption of length $T$ is placed on the posterior as
@@
This objective is implemented in the Hierarchical PU-Net [74] (HPU-Net, depicted in Figure 9). Simply stated, the architecture learns a latent representation at multiple decomposition levels of the U-Net. Furthermore, residual connections in the convolutional layers are necessary to prevent degeneracy of uninformative latent variables with the KL divergence rapidly approaching zero. For similar reasons, the Generalized ELBO with Constrained Optimization (GECO) objective is employed, which extends Equation (28) to
$\mathcal{L}_{\text{GECO}}=\lambda$ $\cdot(\mathbb{E}_{q_{\theta}(\mathbf{Z}|\mathbf{Y},\mathbf{X})}[\,\log p_{\phi}(\mathbf{Y}|\mathbf{z},\mathbf{X})\,]-\kappa)$ (29)
@@
Hyperparameter $\lambda$ is the Lagrange multiplier update through the Exponential Moving Average of the reconstruction, which is constrained to reach target value $\kappa$, empirically set beforehand to an appropriate value. Finally, online negative hard mining is used to only backpropagate 2% of the worst performing pixels, which are stochastically picked with the Gumbel-SoftMax trick [87, 88]. A similar architecture known as PHiSeg [75] takes a similar approach to the HPU-Net. However, instead of placing the residual connections in the convolutional layers, PHiSeg uses these between the latent vectors across the decompositions. Furthermore, deep supervision at each decomposition scale is used to enforce disentanglement across latent variables.
@@
Early methods for uncertainty quantification in medical imaging primarily utilized 2D slices from three-dimensional (3D) datasets, leading to a potential loss of critical spatial information and subtle nuances often necessary for accurate delineation. This limitation has spurred extensive research into 3D probabilistic segmentation techniques with ELBO-based models, aiming to preserve the integrity of entire 3D structures. Initial works [79, 80] demonstrate that the PU-Net can be adapted by replacing all 2D operations with their 3D variants. Crucially, the fusion of the latent sample with 3D extracted features is achieved through a 3D tiling operation. Viviers et al. [78] additionally augment a Normalizing Flow to the posterior density, as described in Section IV-B2. Further enhancements to the architecture include the implementation of the 3D HPU-Net [77], an updated feature network incorporating the attention mechanisms, a nested decoder, and different reconstruction loss components tailored to specific applications [76].
@@
It can be relevant to model the annotators independently in cases with consistent annotator-segmentation pairs in the dataset. This can be achieved by conditioning the learned densities on the annotator itself [81, 89]. For example, features of a U-Net can be combined with samples from annotator-specific Gaussian-distributed posteriors [89]. Considering the approach from Gao et al. [81], generating a segmentation mask is achieved by first sampling an annotator from a categorical prior distribution $\mathcal{C}(\pi_{k}(\mathbf{x}))$, governed by the image conditional parameters $\pi_{k}(\mathbf{x})$ for the $k$-th annotator. Then, samples are taken from its corresponding prior density as $\mathbf{z}_{k}\sim p_{k}(\mathbf{z}_{k})$ to reconstruct a segmentation through an image-conditional decoder $\mathbf{y}=F(\mathbf{x},\mathbf{z}_{k})$. The parameters $\pi_{k}(\mathbf{z}_{k})$ can also be used to weigh the corresponding predictions to express the uncertainty in the prediction ensemble. Additionally, consistency between the model and ground-truth distribution is enforced through an optimal transport loss between the set of predictions and labels.
IV-B3 Denoising Diffusion Probabilistic Models (DDPMs)
Recent developments in generative modeling have resulted in a family of models known as Denoising Diffusion Probabilistic Models [90, 91, 92]. Such models are especially renowned for their expressive power by being able to encapsulate large and diverse datasets. While several perspectives can be used to introduce the DDPMs, we build upon the earlier discussed HVAE (Section IV-B2) to maintain cohesiveness with the overall manuscript.
Specifically, we introduce three additional modifications to the HVAE [93]. Firstly, the latent dimensionality is set equal to the spatial data dimensions, i.e. $d\,$=$\,D$. As a consequence, redundant notation of $\mathbf{Z}$ is removed and $\mathbf{Y}$ is instead subscripted with $t\in\{1,...,T\}$, indicating the encoding depth, where $\mathbf{Y}_{0}$ is the initial segmentation mask. Secondly, the encoding process (or forward process) is predefined as a linear Gaussian model such that
$q(\mathbf{y}_{T}|\mathbf{y}_{0})=p(\mathbf{y}_{0})\prod^{T}_{t=1}q(\mathbf{y}_{t}|\mathbf{y}_{t-1}),$ (30)
where